OpenStudy (anonymous):

WILL GIVE MEDAL!! Find the exact value of the following. In each case show your work and explain the steps

OpenStudy (anonymous):

sin(17pi/6) tan(13pi/4) sec(11pi/3)

OpenStudy (anonymous):

can you help @polaris_s0i

OpenStudy (polaris_s0i):

I am assuming you aren't just supposed to take the arcsin of it and get the answer. I understand what its asking, I'm just trying to figure out how to formulate it.

OpenStudy (mertsj):

\[\frac{17\pi }{6}= 2 \frac{5}{6}\pi \]

OpenStudy (mertsj):

So, the sin of 17pi/6 is equal to the sin of 5pi/6

OpenStudy (mertsj):